intereste payment=575


GST=125


Dividend received=330


Recovery of loans=615


Salary paid=820


Disinvestment=320


Borrowings=1115


Escheat=480


Find :-


1. Budget deficit


2. Revenue deficit


3. Fiscal deficit


4.primary deficit


1
Expert's answer
2021-01-13T10:23:38-0500

Solution:

1.). Budget deficit = Total expenditures – Total income

Total expenditures = 575 + 615 + 820 = 2,010

Total income = 125 + 330 + 320 = 775

Budget deficit = 2,010 – 775 = 1,235

 

2.). Revenue deficit = Total revenue expenditures – Total revenue receipts

Total revenue expenditures = 575 + 615 + 820 = 2,010

Total revenue receipts = 125 + 330 = 455

Revenue deficit = 2,010 – 455 = 1,555

 

3.). Fiscal deficit = Total expenditure – Total receipts excluding borrowings

Total expenditure = 575 + 615 + 820 = 2,010

Total receipts excluding borrowings = 125 + 330 + 320 + 480 = 1,255

Fiscal deficit = 2,010 - 1,255 = 755  895

 

4.). Primary deficit = Fiscal deficit – Interest payment

Fiscal deficit = 755

Interest payment = 575

Primary deficit = 755 – 575 = 180